An environmental scientist needs to use solid NaOH reduce the acidity of a threatened small lake with a volume of 3.00x10^8 L and pH of 4.8 assuming solid NaOH dissociates what mass of NaOH is needed to get the lake from a pH of 4.8 to 6.5?
1
Expert's answer
2020-06-08T15:24:06-0400

PH = 4.8 .

Concentration of H+ = 10-4.8 M.

Concentration Of H+ = 1.6 × 10-5 M.

So, Concentration Of OH- = 10-14 / 1.6 × 10-5 .

Concentration of OH- = 6.25 × 10-10 M.

When PH = 6.5 .

Concentration of H+ = 3.2 × 10-7 M.

Concentration of OH- = 10-14 / 3.2 × 10-7.

Concentration of OH- = 3.125 × 10-8. M.


So, moles of NaOH required =

(3.12 × 10-8 × 3 × 108 - 6.25 × 10-10 × 3 × 108).


Moles of NaOH Required=

(3.12×3 - 6.25×3×10-2).mole.


Moles of NaOH = (9.36 - 0.1875) mole.

= 9.1725 moles.


Mass of NaOH = (9.1725 × 40 ). gm.

Mass Of NaOH = 366.9 gm. Answer.